Joe is three times as old as Ben. Together their ages equal 104.How old is Joe? How old is Ben?
Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52810)

.
Duplicate.


Answer by addingup(3677)

J = 3B
J+B = 104
3B+B = 104
4B = 104
B = 26 Ben is 26
and the problem says J = 3B:
3(26) = 78 Joe is 78
------------------------------
Check:
Joe + Ben = 104
78 + 26 = 104 Correct